1/1

Irmtraut Böttcher,

Herbstgedanken

Argiope information: The view of this version is restricted. Presumably there are copyright issues.

Author of this German poem is Irmtraut Böttcher (∗1918-06-16).

Arg-330-665 (2005-10-26 08:10:27)

Frau Irmtraut Böttcher (Grabenstr. 35, D-58095 Hagen), die Verfasserin des Gedichtes, hat mir, Manfred Retzlaff, freundlicherweise die Veröffentlichung ihres Gedichtes in dieser meiner Internetz-Gedichtsammlung gestattet.

http://www.poezio.net/version?poem-id=330&version-id=665